# Private Fine-tuning of Large Language Models with Zeroth-order Optimization

### v1

@@ 5 DISCUSSION
In Section 4.2 we showed that DP-ZO obtains superior results to DP-SGD by scaling to larger models. Now we examine the amount of engineering effort necessary to scale DP-SGD to larger models, a topic on which many papers have been written (Bu et al., 2023a, b; Yousefpour et al., 2021; Li et al., 2022b; He et al., 2023). We find that DP-ZO seamlessly scales to larger models and believe its simplicity presents a compelling alternative to DP-SGD for practitioners.
DP-SGD.
@@
Differentially Private Stochastic Gradient Descent (DP-SGD) (Song et al., 2013; Abadi et al., 2016) is the standard privacy-preserving training algorithm for training neural networks on private data, with an update rule given by $w^{(t+1)}=w^{(t)}-\frac{\eta_{t}}{|B|}\left(\sum_{i\in B}\frac{1}{\text{c}}\texttt{clip}_{\text{c}}(\nabla\ell(x_{i},w^{(t)}))+\sigma\xi\right)$ where the changes to SGD are the per-sample gradient clipping $\texttt{clip}_{\text{c}}(\nabla\ell(x_{i},w^{(t)}))=\frac{c\times\nabla\ell(x_{i},w^{(t)})}{\max(c,||\nabla\ell(x_{i},w^{(t)})||_{2})}$ and addition of noise sampled from a $d$-dimensional Gaussian distribution $\xi\sim\mathcal{N}(0,1)$ with standard deviation $\sigma$. DP-SGD is the marquee algorithm for privacy-preserving machine learning, but it requires implementing per-example gradient clipping. This creates a slew of challenges for deploying DP-SGD.
Computational and Memory Challenges in DP-SGD. DP-SGD requires the computation of per-example gradients, which can be naively implemented by storing each gradient in the batch separately. This approach inflates the memory overhead by a factor of $B$, where $B$ is the batch size. Tensorflow Privacy avoids this issue by clipping microbatches rather minibatches, which does not slow down training but increases the noise added and therefore hurts utility. Jax can automatically vectorize the per-sample gradient computation, but training is still slowed down. Recently, specialized libraries have been developed that instead analytically compute the norm of the gradients for different layers (Li et al., 2022b; Bu et al., 2023c). This requires actually implementing the computation, which is challenging for new layers. If a practitioner wants to train LLaMA2 and efficiently compute the per-example gradient norms, they would first have to derive the analytical formula for the norm of GroupedQueryAttention (Touvron et al., 2023), which can represent a nontrivial amount of engineering. This complexity is further compounded when considering models with parameter sharing, such as the OPT networks we use in our experiments, as this is generally not compatible with analytical norm clipping methods (lxuechen, 2022). Although network architectures exist that do not require parameter sharing, the best pretrained models (Zhang et al., 2022; Touvron et al., 2023) use parameter sharing because it makes the model use the same representation for predicting tokens at the next step as well as for decoding.
@@ 5.1 DP-ZO Scales Seamlessly
In order to train models as large as OPT-66B, whose parameters cannot be loaded into memory on a single A100 80GB GPU, we need to implement some form of parallelism across GPUs. We now discuss how easy this is for DP-ZO (in the simplest form, just running DP-ZO on a machine with 2 GPUs will prompt HuggingFace to implement naive model parallelism) and how challenging it can be for approaches that require per-example gradient clipping such as DP-SGD.
Data Parallelism in DP-ZO. To synchronize model state between GPUs in data-parallel-DP-ZO, we just transfer the random seed and its corresponding fp16 scalar step size; this is just a few bytes. However, first-order approaches such as DP-SGD require the transfer of gradients across devices to update all the models, necessitating expensive allgather and reduce operations. This communication overhead is $1.5d$ in PyTorch FSDP, where $d$ is the size of the model.
DP-ZO Does Not Store Gradients. DP-ZO does not store activations or gradients in the forward pass, thus conserving memory. DP-SGD needs to store the activations at each GPU under pipeline parallelism in order to clip the per-example gradient (He et al., 2023), which will fill up the GPU memory and limit new microbatches from being processed.
DP-ZO Easily Integrates With Mixed Precision Training. All of our DP-ZO experiments use mixed precision; we load the model and compute the loss in half precision. Mixed precision training is much faster than full precision and also has a smaller memory footprint; these factors contribute to DP-ZO’s ability to train much larger models. By contrast, implementing DP-SGD with half precision is challenging. Yu et al. (2022) do not modify their implementation of DP-SGD to support half precision and report worse performance as a result. Li et al. (2022b) detail an algorithm that interleaves loss scaling with gradient clipping to avoid underflow while maintaining utility. This implementation is nontrivial and as of today is not available in Opacus (ffuuugor, 2022).
@@ 6 RELATED WORK
Recent studies have leveraged DP-SGD to fine-tune large language models. Li et al. (2022b) provide methods for fine-tuning large language models with DP-SGD by ghost clipping to mitigate the memory burden of per-sample gradient clipping. Yu et al. (2022) report compelling results by only updating a sparse subset of the LLMs with parameter efficient fine-tuning (PEFT) methods such as LoRA (Hu et al., 2022). He et al. (2023) leverage group-wise clipping with adaptive clipping threshold and privately fine-tune the 175 billion-parameter GPT-3. Duan et al. (2023); Li et al. (2022b) also consider private prompt tuning by adding noise to the soft prompt (Li and Liang, 2021; Lester et al., 2021). Du et al. (2023) add non-i.i.d. noise from a matrix Gaussian distribution to directly perturb embedding in the forward pass of language models. With the emergence in-context learning of large language models (Brown et al., 2020), recent works (Duan et al., 2023; Wu et al., 2023; Tang et al., 2023b) study private in-context learning of large language models without fine-tuning.
Private Zeroth-order Optimization.
@@
Most recently, a concurrent work (Zhang et al., 2023) also considers the same DP-SPSA algorithm for zeroth-order optimization. Our method and Zhang et al. (2023) are functionally the same up to constants, and our work focuses on an empirical evaluation of the method, whereas Zhang et al. (2023) extends the convergence analysis of Malladi et al. (2023) to DP as shown in Appendix B of Zhang et al. (2023). There is a slight difference for the generation of random perturbation of Zhang et al. (2023) and our Algorithm 1. Zhang et al. (2023) uses the random unit vector for the perturbation and the convergence analysis is based on such set-up, whereas our perturbation is a normally distributed vector. Note that Algorithm 1 and 2 in Zhang et al. (2023) also scales the unit vector by the square root of the model dimension, so the two approaches are functionally the same. We reimplemented our perturbation method based on the algorithms in Zhang et al. (2023), and we obtain the same performance of $81.3\%$ F1 with $(1,10^{-5})$-DP for OPT-13B by DP-ZO by LoRA finetuning on SQuAD.
@@
Gratton et al. (2022) analyze the intrinsic privacy of the zeroth-order optimization for DP-ADMM (Huang et al., 2020) in distributed learning.  Gratton et al. (2022) state that if the output of the zeroth-order method itself follows Gaussian distribution, the noise can be analyzed as the Gaussian mechanism and provide intrinsic privacy. However, this is merely stated as an assumption for lemma 1. To the best of our knowledge there is no work that proves that the zeroth-order gradient estimator can actually be analyzed as the sum of an unbiased gradient estimator and some Gaussian error term.
@@ 7 CONCLUSION
DP-SGD has been the de-facto private training method of the last decade. In this work we propose DP-ZO, a novel method for private fine-tuning that privatizes the zeroth-order update by adding noise to the difference in loss between two perturbations. DP-ZO’s unique univariate privatization unlocks training larger models with better parallelism than DP-SGD. In particular, DP-ZO can train OPT-66B at $(1,10^{-5})$-DP in the challenging few-shot setting with only $1.86\%$ performance degradation due to privacy, beating DP-SGD by $1.71\%$. We also achieve the first non-trivial pure $\varepsilon$-DP result by instantiating DP-ZO with the Laplace mechanism. We anticipate that future work can further study these design choices, integrate more DP mechanisms into DP-ZO, and apply it to the vision domain.
Acknowledgement
This work was supported in part by the National Science Foundation under grant CNS-2131938, the ARL’s Army Artificial Intelligence Innovation Institute (A2I2), Schmidt DataX award, and Princeton E-ffiliates Award.
@@ Appendix A Privacy Analysis
Proposition A.1 (Basic Composition theorem (Dwork and Roth, 2014)).
@@
Proposition A.2 (Privacy Amplification via Subsampling (Balle et al., 2018)).
@@
As outlined in Line 11, the $\ell_{2}$ sensitivity of Alg. 1 is $\mathrm{C}$ and we are adding $\mathcal{N}(0,\mathrm{C}^{2}\sigma^{2})$ noise to the estimated loss. We analyze the composition of subsampled Gaussians with the privacy loss variable accounting approach of Gopi et al. (2021).
@@
We first analyze the privacy cost for one step by the Laplace mechanism. At each step, we sample a new batch of data with the sample rate of $p=B/|\mathcal{D}|$. As outlined in 11, the $\ell_{1}$ sensitivity of Algorithm 1 is $\mathrm{C}$. By Proposition 2.3, the privacy cost at one step would cost $(1/\sigma,0)$-DP on this batch. By Proposition A.2, the privacy cost at one step would cost $(\log(1+p\cdot(e^{1/\sigma}-1)),0)$-DP on the full dataset $\mathcal{D}$. By  Proposition A.2, the privacy cost of Algorithm 1 instantiated with Laplace mechanism satisfies $(T\cdot\log(1+p\cdot(e^{1/\sigma}-1)),0)$-DP. We provide the privacy parameters we used for pure $\varepsilon$-DP by the Laplace mechanism in Table 9.
We can also get tighter composition of $\varepsilon$ with relaxation to $\delta>0$. Specifically, we can instantiate the PRV of random response with $(\log(1+p\cdot(e^{1/\sigma}-1)),0)$ because the dominating pair for random response is a dominating pair for the pure DP mechanism. Then, we can use the numerical composition of Pure DP PRV by Gopi et al. (2021). We use this privacy accounting method for the approximate $(\varepsilon,\delta)$-DP results for Laplace mechanisms in the main paper.
In addition, we could also compute the privacy cost of composition for the Laplace mechanism by Monte Carlo based DP accountant (Wang et al., 2023). Note that since we are dealing with scalar values, our mechanism in each iteration will be a one dimensional Laplace mechanism. Let $b$ be the scale of Laplace noise, $p$ the sub-sampling rate, and assume the sensitivity is $1$, and assume we are doing composition for $T$ iterations, each iteration with sampling rate $p$. By Zhu et al. (2022) we know that the pair of distribution $(P,Q)$ dominating pair for a single dimensional Laplace mechanism, where $P$ and $Q$ are distributed according to the following pdfs,
@@
Therefore, ($P,(1-p)\cdot P+p\cdot Q$) is the dominating pair for the sub-sampled Laplace. We plug this into the standard Monte-Carlo accountant of Wang et al. (2023) (without importance sampling, see Algorithm 2 and Theorem 10 in Wang et al. (2023)) while using $10^{10}$ samples to calculate the $\delta$ at a given value of $\varepsilon$. Also, using the analytical accountant explained above, we always make sure that $\mathbb{E}[\hat{\delta}_{MC}^{2}]$ is bounded by $10^{-8}$ (We use the fact that $\mathbb{E}[\hat{\delta}_{MC}^{2}]$ is bounded by $\mathbb{E}[PRV^{2}]$ and the fact the $PRV$ is always bounded for Laplace mechanism.). This ensures that the error of our estimation of $\delta$ is at most $10^{-8}$ with probability at least $1-10^{-5}$. Putting all together, for all reported values of $\varepsilon$, our $\delta$ is bounded by $10^{-5}$, with probability at least $0.99999$. This privacy analysis is tighter with $\varepsilon$ is high compared to the former analysis which use the pure DP PRV accountant. This is consistent with the intuition. As we increase the distance between the Laplace dominating pairs, the probability of sampling points from the area between the centers increases. And that is where the Laplace Mechanism is different from the Randomized Response. We present the accounting results for the Laplace method to achieve $(\varepsilon,\delta)$-DP by these two accounting methods in Table 9. Table 9 shows that the Monte Carlo based DP accountant can give tighter analysis for the Laplace mechanism for $(\varepsilon,\delta)$-DP than the pure $\varepsilon$-DP PRV method.
Table 8: $\varepsilon$-DP by Laplace. BSZ=16, Steps=10000.
Table 9: $(\varepsilon,\delta)$-DP guarantee for Laplace. $\delta=10^{-5}$. $|\mathcal{D}|=1000$. BSZ=16, Steps=10000.
@@
16.5 0.33 0.35
11.76 0.48 0.5
6.5 0.9 1
2.19 $2.9$ 4
Table 9: $(\varepsilon,\delta)$-DP guarantee for Laplace. $\delta=10^{-5}$. $|\mathcal{D}|=1000$. BSZ=16, Steps=10000.

### v2

@@ 5 DISCUSSION
In Section 4.2 we showed that DP-ZO obtains competitive privacy-utility tradeoff. Now we examine the amount of engineering effort necessary to scale DP-SGD to larger models, a topic on which many papers have been written (Bu et al., 2023b, c; Yousefpour et al., 2021; Li et al., 2022b; He et al., 2023; Bu et al., 2023a). We find that DP-ZO seamlessly scales to larger models and believe its simplicity presents a compelling alternative to DP-SGD for practitioners.
@@
DP-SGD. Differentially Private Stochastic Gradient Descent (DP-SGD) (Song et al., 2013; Abadi et al., 2016) is the standard privacy-preserving algorithm to train models on private data, with an update rule given by $w^{(t+1)}=w^{(t)}-\frac{\eta_{t}}{|B|}\left(\sum_{i\in B}\frac{1}{\text{c}}\texttt{clip}_{\text{c}}(\nabla\ell(x_{i},w^{(t)}))+\sigma\xi\right)$ where the changes to SGD are the per-sample gradient clipping $\texttt{clip}_{\text{c}}(\nabla\ell(x_{i},w^{(t)}))=\frac{c\times\nabla\ell(x_{i},w^{(t)})}{\max(c,||\nabla\ell(x_{i},w^{(t)})||_{2})}$ and addition of noise sampled from a $d$-dimensional Gaussian distribution $\xi\sim\mathcal{N}(0,1)$ with standard deviation $\sigma$. DP-SGD is the marquee algorithm for privacy-preserving machine learning, but it requires implementing per-example gradient clipping. This creates a slew of challenges for deploying DP-SGD.
Computational and memory challenges in DP-SGD. DP-SGD requires the computation of per-example gradients, which can be naively implemented by storing each gradient in the batch separately. This approach inflates the memory overhead by a factor of $B$, where $B$ is the batch size. Tensorflow Privacy avoids this issue by clipping microbatches rather minibatches, which does not slow down training but increases the noise added and therefore hurts utility. Jax can automatically vectorize the per-sample gradient computation, but training is still slowed down. Recently, specialized libraries have been developed that instead analytically compute the norm of the gradients for different layers (Li et al., 2022b; Bu et al., 2023d; Ding et al., 2024). This requires actually implementing the computation, which is challenging for new layers. Parameter efficient fine-tuning methods (Yu et al., 2022; Bu et al., 2024) can help reduce such computation cost by reducing the number of trainable parameters. However, as discussed in Section 4.3, those methods still incur much more memory cost than DP-ZO for long sequences. Besides, when model cannot be loaded into a single GPU, model parallelism is needed to load large models across several GPUs and per-example gradient norm clipping requires additional implementation, both for gradient clipping and communication across device. He et al. (2023) investigate how to make group-wise gradient clipping efficient and achieve good performance in DP-SGD and fine-tunes 175B GPT3 model with 16 V100 GPUs each with 32 gigabytes of VRAM. Bu et al. (2023a) implements DP-SGD based on Zero Redundancy Optimizer (Rajbhandari et al., 2020) to scale model size up to GPT-100B and maintain efficiency. It currently also only supports layer-wise or block-wise clipping.
@@ 5.1 DP-ZO Scales Seamlessly
We now discuss the advantage of DP-ZO that can scale to large language models seamlessly. Note that DP-ZO inherits the seamless scalability from ZO as a result of only additional computation cost on loss. DP-ZO achieves comparable performance as DP-SGD. In contrast, DP-SGD incurs more computational and memory challenges than SGD due to per-example gradient clipping.
Model parallelism in DP-ZO. To train large models like OPT-66B, whose parameters cannot be loaded into memory on a single A100 GPU, we need to implement some form of parallelism across GPUs. It is easy for such parallelism in DP-ZO (in the simplest form, just running DP-ZO on a machine with 2 GPUs will prompt HuggingFace to implement naive model parallelism), while much more effort in DP-SGD as we previously discussed.
Data parallelism in DP-ZO. To synchronize model state between GPUs in data-parallel-DP-ZO, we just transfer the random seed and its corresponding half-precision float16 scalar step size; this is just a few bytes. However, first-order approaches such as DP-SGD require the transfer of gradients across devices to update all the models, necessitating expensive allgather and reduce operations. This communication overhead is $1.5d$ in PyTorch FSDP, where $d$ is the size of the model.
DP-ZO does not store gradients. DP-ZO does not store activations or gradients in the forward pass, thus conserving memory. DP-SGD needs to store the activations at each GPU under pipeline parallelism to clip the per-example gradient (He et al., 2023), which will fill up the GPU memory and limit new microbatches from being processed.
@@ 6 RELATED WORK
Privacy preserving large language models. Recent studies have leveraged DP-SGD to fine-tune large language models. Li et al. (2022b) provide methods for fine-tuning large language models with DP-SGD by ghost clipping to mitigate the memory burden of per-sample gradient clipping. Yu et al. (2022) report compelling results by only updating a sparse subset of the LLMs with parameter efficient fine-tuning (PEFT) methods such as LoRA (Hu et al., 2022). He et al. (2023) leverage group-wise clipping with adaptive clipping threshold and privately fine-tune the 175 billion-parameter GPT-3. Duan et al. (2023); Li et al. (2022b) also consider private prompt tuning by adding noise to the soft prompt (Li and Liang, 2021; Lester et al., 2021). Du et al. (2023) add non-i.i.d. noise from a matrix Gaussian distribution to directly perturb embedding in the forward pass of language models. With the emergence in-context learning of large language models (Brown et al., 2020), recent works (Duan et al., 2023; Wu et al., 2024; Tang et al., 2023b) study private in-context learning of large language models without fine-tuning.
@@
Private zeroth-order optimization. Most recently, a concurrent work (Zhang et al., 2024a) also considers the same DP-SPSA algorithm for zeroth-order optimization. Our method and Zhang et al. (2024a) are functionally the same up to constants, and our work focuses on an empirical evaluation of the method, whereas Zhang et al. (2024a) extends the convergence analysis of Malladi et al. (2023) to DP as shown in Appendix B of Zhang et al. (2024a). There is a slight difference for the generation of random perturbation of Zhang et al. (2024a) and our Algorithm 1. Zhang et al. (2024a) uses the random unit vector for the perturbation and the convergence analysis is based on such set-up, whereas our perturbation is a normally distributed vector. Note that Algorithm 1 and 2 in Zhang et al. (2024a) also scales the unit vector by the square root of the model dimension, so the two approaches are functionally the same. We reimplemented our perturbation method based on the algorithms in Zhang et al. (2024a), and we obtain $82.32_{0.82}$ for fine-tuning OPT-13B by LoRA finetuning on SQuAD with $(1,10^{-5})$-DP, that is comparable as our main result in Table 1. Besides, Our work provides a systematic study of DP-ZO and DP-SGD on privacy-utility trade-off including different models and different noise mechanisms, as well as computation cost and privacy implication.
Zhang et al. (2024b) study private zeroth-order nonsmooth nonconvex optimization. Their work incorporates two zeroth-order estimators to reduce variance and samples $d$ (model dimension) i.i.d. estimators for each data point to achieve optimal dimension dependence. Zhang et al. (2024b) leverage the tree mechanism (Dwork et al., 2010; Chan et al., 2011) on disjoint data to ensure the privacy cost of the algorithm. The main focus of our work is private fine-tuning of large language models and one estimator for each batch could successfully converge in this set-up. Therefore, we only need to privatize such scalar. We leave the investigation on the private zeroth-order for more than one estimators such as the variance reduction method proposed in Zhang et al. (2024b) as future work.
@@
Gratton et al. (2022) analyze the intrinsic privacy of the zeroth-order optimization for DP-ADMM (Huang et al., 2020) in distributed learning. Their work states that if the output of the zeroth-order method itself follows Gaussian distribution, the noise can be analyzed as the Gaussian mechanism and provide intrinsic privacy. However, this is merely stated as an assumption for lemma 1. To the best of our knowledge there is no work that proves that the zeroth-order gradient estimator can actually be analyzed as the sum of an unbiased gradient estimator and some Gaussian error term.
@@ 7 CONCLUSION
DP-SGD has been the de-facto private training method of the last decade. In this work we propose DP-ZO, a novel method for private fine-tuning that privatizes the zeroth-order update by adding noise to the difference in loss between two perturbations. DP-ZO’s unique univariate privatization unlocks training larger models with better parallelism than DP-SGD. DP-ZO provides a strong privacy-utility trade-off across different tasks, model sizes, dataset sizes, and DP mechanisms. We anticipate that future work can further study these design choices, integrate more DP mechanisms into DP-ZO, and apply it to the vision domain.
@@ Appendix A Privacy Analysis
Proposition A.1 (Basic Composition theorem [Dwork and Roth, 2014]).
@@
Proposition A.2 (Privacy Amplification via Subsampling [Balle et al., 2018]).
@@
As outlined in Line 11, the $\ell_{2}$ sensitivity of  Algorithm 1 is $\mathrm{C}$ and we are adding $\mathcal{N}(0,\mathrm{C}^{2}\sigma^{2})$ noise to the estimated loss. We analyze the composition of subsampled Gaussians with the privacy loss variable accounting approach of Gopi et al. [2021].
@@
Laplace mechanism can gives a pure DP guarantee of $\delta=0$ which can be of interest in some scenarios. Here we first analyze the pure $\varepsilon$-DP guarantee provided by laplace mechanism and then provide the analysis for approximate $(\varepsilon,\delta)$-DP analysis.
Pure $\varepsilon$-DP by Laplace mechanism. We use data in a single batch instead of all training data to compute the gradients in each updates. For the privacy analysis for Laplace mechanism in Algorithm 1, when we sample each batch in the poisson manner, we could leverage Proposition A.2 to compute the private amplification by subsampling. We first analyze the privacy cost for one step by the Laplace mechanism. At each step, we sample a new batch of data with the sample rate of $p=B/|\mathcal{D}|$. As outlined in 11, the $\ell_{1}$ sensitivity of Algorithm 1 is $\mathrm{C}$. By Proposition 2.3, the privacy cost at one step would cost $(1/\sigma,0)$-DP on this batch. By Proposition A.2, the privacy cost at one step would cost $(\log(1+p\cdot(e^{1/\sigma}-1)),0)$-DP on the full dataset $\mathcal{D}$. By Proposition A.1, the privacy cost of Algorithm 1 instantiated with Laplace mechanism satisfies $(T\cdot\log(1+p\cdot(e^{1/\sigma}-1)),0)$-DP.
We provide the privacy parameters we used for pure $\varepsilon$-DP by the Laplace mechanism in Table 12.
Approximate $\varepsilon$-DP by Laplace mechanism. We can also get tighter composition of $\varepsilon$ with relaxation to $\delta>0$. The most straight forward way is to instantiate the PRV of random response with $(\log(1+p\cdot(e^{1/\sigma}-1)),0)$ because the dominating pair for random response is a dominating pair for the pure DP mechanism. Then, we can use the numerical composition of Pure DP PRV by Gopi et al. [2021]. Note that this method is agnostic to the DP mechanisms used for pure $\varepsilon$-DP. We now provide a more fine-grained privacy analysis for the laplace mechanism. Specifically, we could compute the privacy cost of composition for the Laplace mechanism by Monte Carlo based DP accountant [Wang et al., 2023]. Note that since we are dealing with scalar values, our mechanism in each iteration will be a one dimensional Laplace mechanism. Let $b$ be the scale of Laplace noise, $p$ the sub-sampling rate, and assume the sensitivity is $1$, and assume we are doing composition for $T$ iterations, each iteration with sampling rate $p$. By Zhu et al. [2022] we know that the pair of distribution $(P,Q)$ dominating pair for a single dimensional Laplace mechanism, where $P$ and $Q$ are distributed according to the following pdfs,
@@
Therefore, ($P,(1-p)\cdot P+p\cdot Q$) is the dominating pair for the sub-sampled Laplace. We plug this into the standard Monte-Carlo accountant of Wang et al. [2023] (without importance sampling, see Algorithm 2 and Theorem 10 in Wang et al. [2023]) while using $10^{10}$ samples to calculate the $\delta$ at a given value of $\epsilon$. Also, using the analytical accountant explained above, we always make sure that $\mathbb{E}[\hat{\delta}_{MC}^{2}]$ is bounded by $10^{-8}$ (We use the fact that $\mathbb{E}[\hat{\delta}_{MC}^{2}]$ is bounded by $\mathbb{E}[PRV^{2}]$ and the fact the $PRV$ is always bounded for Laplace mechanism.). This ensures that the error of our estimation of $\delta$ is at most $10^{-8}$ with probability at least $1-10^{-5}$. Putting all together, for all reported values of $\epsilon$, our $\delta$ is bounded by $10^{-5}$, with probability at least $0.99999$. This privacy analysis is tighter with $\varepsilon$ is high compared to the former analysis which uses the pure DP PRV accountant. This is consistent with the intuition. As we increase the distance between the Laplace dominating pairs, the probability of sampling points from the area between the centers increases. And that is where the Laplace Mechanism is different from the Randomized Response. We present the accounting results for the Laplace method to achieve $(\varepsilon,\delta)$-DP by these two accounting methods in Table 12. Table 12 shows that the Monte Carlo based DP accountant can give tighter analysis for the Laplace mechanism for $(\varepsilon,\delta)$-DP than the pure $\varepsilon$-DP PRV method.
Table 11: $\varepsilon$-DP by Laplace. BSZ=20, Steps=2000.
@@
Table 12: $(\varepsilon,\delta)$-DP guarantee for Laplace. $\delta=10^{-5}$. $|\mathcal{D}|=1000$. BSZ=16, Steps=75000.
@@
30.8 0.5 0.51
16.3 1 1.04
4.6 4 4.70